OpenStudy (anonymous):

Miley’s dog, Ella loves to play outside. Ella’s 15 ft leash is tied to a stake in the ground. How much area does she have to play? A. 807.5 ft^2 B. 78.5 ft^2 C. 47.1 ft^2 D. 706.9 ft^2

OpenStudy (anonymous):

D. because, pi * 15^2 = 225 * pi =~ 706.858 ft^2

OpenStudy (anonymous):

hello? did that help.. you even check for yourself I gave you the right steps lol